Identify the following compound:
C5H10O

NMR: δ 9.8 (1 H, s), δ 1.1 (9 H, s)
Answer #1
Concepts and reason

This problem is based on the concept of NMR spectroscopy.

NMR (nuclear magnetic resonance) spectroscopy is a well-known technique in organic chemistry which is used to determine the structure of the organic compound. It works on the principle in the presence of an external magnetic field some nuclei which have spin are transferred from lower energy level to higher energy.

Fundamentals

The double bond equivalence/ degree of unsaturation gives an idea about the number of double bonds or rings present. It can be calculated as:

DBE=CH2N2+1{\rm{DBE = C - }}\frac{{\rm{H}}}{{\rm{2}}} - \frac{{\rm{N}}}{2} + 1

Here, C, H and N is number of carbon, hydrogen and nitrogen atom respectively. Determination of the DBE, gives us a preliminary idea about the structure.

By substituting the value of C, H and N as 5,10 and 0 in equation (1), we calculate the Double bond equivalence as:

DBE=510202+1=1\begin{array}{c}\\{\rm{DBE = 5 - }}\frac{{10}}{{\rm{2}}} - \frac{0}{2} + 1\\\\ = 1\\\end{array}

By analyzing, the given values of the shift, the structure is found to be:

HC=0
H3C-
CH3
CH3

Ans:

The structure of the compound is as follows:

HC=0
H3C-
CH3
CH3